OpenStudy (anonymous):

1.) Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 2x-y=2 3x-2y=11 2.)Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) x+y=17 y=x+7

OpenStudy (anonymous):

so for example, why dont you try to make it easy and on number 1, the top equation, try to plug in 0 for x, and solve for y. or you can try 1, 2 or 3 and solve for the other variable.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

First, solve the first equation for a variable, so you get either x=something or y=something. Then, in the second equation, plug in for x or y (whichever one you solved for) what you got in the first equation. Then, solve for the only variable left.

OpenStudy (anonymous):

I'll do the first one! 2x-y=2 3x-2y=11 So we would isolate y; so y=2x-1 Then plug in y=2x-1 into 3x-2y=11 So 3x-2(2x-1)=11 Then solve. 3x-4x+2=11 -x=9 So x=-9 Then plug that into the original So 2(-9)-y=2 -18-y=2 y=-20 (-9,-20)

OpenStudy (anonymous):

thanks for the help